I would please like your help with this problem: 5x-3(2x-1)=5-3(-x+2)
Answer by fractalier(6550)

Okay, from
5x - 3(2x - 1) = 5 - 3(-x + 2)
let us expand everything, then collect terms...
5x - 6x + 3 = 5 + 3x - 6
-x + 3 = 3x - 1
Now addx to both sides and then add 1 to both sides.
3 = 4x - 1
4 = 4x
Now divide by 4 and we get
x = 1